OpenStudy (anonymous):

26=6(5-4f) help plzz.

jimthompson5910 (jim_thompson5910):

26=6(5-4f) 26=30-24f 26-30=-24f -4 = -24f -4/(-24) = f 1/6 = f f = 1/6

OpenStudy (anonymous):

26 = 30 -24 f -4 = -24 f 1/6 = f f = 1/6